Determine whether the given ordered pairs are solutions of the equation x-6y=-44
A.(5,0) B.(-8,6)

Answer by ewatrrr(24785)


Hi,
x - 6y = -44
substituting (5,0)
5 - 0 = -44 Ordered pair (5,0)not a solution
substituting (-8,6)
-8 - 6*6 = -44
-8 - 36 = -44 Ordered pair (-8,6) is a solution